# Is GW190521 a gravitational wave echo of wormhole remnant from another universe?

### v1

@@ III.2 Bayesian comparison with BBHs model
It is also necessary to assess the significance of our echo-for-wormhole hypothesis relative to the BBHs hypothesis. The Bayesian model comparison Christensen:2022bxb can be performed by computing the logarithmic Bayes factor. The log Bayes factor between the different source hypothesis ($\mathcal{H}_{i}$) and the noise-only hypothesis ($\mathcal{H}_{0}$) is $\mathrm{ln}\mathcal{B}=\mathrm{ln}{p(d|\mathcal{H}_{i},I)\over p(d|\mathcal{H}_{0},I)}$, where $p(d|\mathcal{H}_{i},I)$ represents the evidence for hypothesis $\mathcal{H}_{i}$. By using the nested sampling package dynesty, we have
$\mathrm{ln}\mathcal{B}^{\text{Echo}}_{\text{BBH}}\simeq-2.9.$ (10)
The Bayesian model comparison slightly favors the BBHs hypothesis, however, this value of log Bayes factor is not significant enough in supporting the BBHs hypothesis again our echo-for-wormhole hypothesis. Thus our model might be an alternative interpretation for the GW190521 event.
IV Discussion
@@
In this paper, we investigate the possibility of interpreting GW190521 as an echo produced by a postmerger wormhole, created from the inspiral and merger of BBHs in another universe and connecting it to our own. Our result indicates that our echo-for-wormhole model yields a network SNR comparable to that of the BBHs model reported by LIGO-Virgo collaboration, thus our model might be a viable and informative hypothesis for the GW190521 event.
@@
It is usually thought that the ringdown signals of wormhole consist of a series of continuous echo pulses, here we consider GW190521 as the first pulse and other undetectable assuming that the wormhole rapidly pinches off and eventually collapses into a black hole Wang:2018mlp or the amplitudes of subsequent echoes fall below the sensitivity threshold of current detectors, e.g.Abedi:2021tti . The issues on the physical reliability of our model are interesting for further studying. The Morris-Thorne wormhole requires the matter with negative energy around the wormhole throat Poisson:1995sv , which might be related to current cosmological observations, e.g.Ye:2020btb ; Wang:2024hwd . Recent other possible observations for wormholes have been explored e.g.Chen:2022tog ; Chen:2024tss . The current Bayesian results indicate that our model is not superior to the BBHs model, however, it might be expected that future enhancements to the echo-for-wormhole waveform templates, incorporating more detailed physical characteristics, hold the potential to significantly improve the corresponding Bayesian factor.
@@
How to accurately identify more such GW events will be significant if our echo-for-wormhole model represents a potential hypothetical source for the GW190521 event. It is noted that recently the LIGO-Virgo collaboration has reported the detection of GW231123 on November 23, 2023, an event sharing a similar burst-like short duration nature with GW190521 LIGOScientific:2025rsn , this event has been motivating intensive investigation on the nature of its source, e.g.DeLuca:2025fln ; Li:2025fnf ; Cuceu:2025fzi ; Delfavero:2025lup ; Gottlieb:2025ugy ; Tanikawa:2025fxw . Therefore, a systematic model comparison test including various possible sources for such short duration GW signals might be necessary for better understanding the physical origins of corresponding GW events.
@@ Acknowledgments
Appendix A Sky locations in our analysis
In the initial analysis, we adopted the same sampling configuration as in the main text but did not impose any constraints on the sky location for either model. The parameter estimation results, summarized in Tab. 3, show that both models yield sky positions broadly consistent with that of the ZTF19abanrhr flare Graham:2020gwr . To mitigate the impact of source localization uncertainty on Bayesian model selection, we fixed the sky position for both the echo-for-wormhole and BBH models to (RA, DEC)=(3.35847,0.60781)=(3.35847,~0.60781), and computed the Bayes factors before (GW) and after (GW+EM) applying this constraint. The results show that although fixing the sky position increases the Bayes factors for both models, it does not produce a significant change in their relative comparison.
Source Right Ascension [rad] Declination [rad] $\mathrm{ln}\mathcal{B}^{\mathrm{GW+EM}}_{\mathrm{GW}}$
ZTF19abanrhr $3.35847$ $0.60781$
BBHs $3.35^{+2.64}_{-0.14}$ $0.45^{+0.19}_{-1.34}$ 3.44
Echo $3.52^{+0.60}_{-0.28}$ $0.62^{+0.28}_{-1.38}$ 3.21

### v2

@@ 3.2 Bayesian comparison with BBHs model
It is also necessary to assess the significance of our echo-for-wormhole hypothesis relative to the BBHs hypothesis. The Bayesian model comparison [47] can be performed by computing the logarithmic Bayes factor. The log Bayes factor between the different source hypothesis ($\mathcal{H}_{i}$) and the noise-only hypothesis ($\mathcal{H}_{0}$) is $\mathrm{ln}\mathcal{B}=\mathrm{ln}{p(d|\mathcal{H}_{i},I)\over p(d|\mathcal{H}_{0},I)}$, where $p(d|\mathcal{H}_{i},I)$ represents the evidence for hypothesis $\mathcal{H}_{i}$. By using the nested sampling package dynesty, we have
$\mathrm{ln}\mathcal{B}^{\text{Echo}}_{\text{BBH}}\simeq-2.9.$ (3.3)
This negative log Bayes factor indicates that the BBHs hypothesis is favored over our echo-for-wormhole hypothesis by the data. Given the simplified nature of our current echo modeling—namely a nonspinning setup and a first-pulse template without an explicit preceding ringdown—a more complete treatment including rotation and a self-consistent modeling of possible later echoes will be required to fully assess this scenario.
4 Discussion
@@
In this paper, we investigate the possibility of interpreting GW190521 as an echo produced by a postmerger wormhole, created from the inspiral and merger of BBHs in another universe and connecting it to our own. Our result indicates that our echo-for-wormhole model yields a network SNR comparable to that of the BBHs model reported by the LVK collaboration, and our analysis provides a proof-of-principle framework to test such an interpretation for the GW190521 event.
@@
It is usually thought that the ringdown signals of wormhole consist of a series of continuous echo pulses, here we consider GW190521 as the first pulse and other undetectable assuming that the wormhole rapidly pinches off and eventually collapses into a black hole [31] or the amplitudes of subsequent echoes fall below the sensitivity threshold of current detectors, e.g.[18]. The issues on the physical reliability of our model are interesting for further studying. The Morris-Thorne wormhole requires the matter with negative energy around the wormhole throat [22], which might be related to current cosmological observations, e.g.[48, 49]. Recent other possible observations for wormholes have been explored e.g.[50, 51]. The current Bayesian results indicate that the BBH hypothesis is favored over our model, and it might be expected that future enhancements to the echo-for-wormhole waveform templates, incorporating more detailed physical characteristics, will be important to assess the robustness of the Bayes factor.
@@
Although the BBH hypothesis is favored for GW190521, how to accurately identify more such short-duration GW events and systematically test alternative hypotheses will be significant. It is noted that recently the LVK collaboration has reported the detection of GW231123 on November 23, 2023, an event sharing a similar burst-like short duration nature with GW190521 [52], this event has been motivating intensive investigation on the nature of its source, e.g.[53, 54, 55, 56, 57, 58]. Therefore, a systematic model-comparison analysis including various possible sources for such short duration GW signals might be necessary for better understanding the physical origins of corresponding GW events.
Appendix A Sky locations in our analysis
In the initial analysis, we adopted the same sampling configuration as in the main text but did not impose any constraints on the sky location for either model. The parameter estimation results, summarized in Tab. 3, show that both models yield sky positions broadly consistent with that of the ZTF19abanrhr flare [79]. To mitigate the impact of source localization uncertainty on Bayesian model selection, we fixed the sky position for both the echo-for-wormhole and BBH models to (RA, DEC)=(3.35847,0.60781)=(3.35847,~0.60781), and computed the Bayes factors before (GW) and after (GW+EM) applying this constraint. The results show that although fixing the sky position increases the Bayes factors for both models, it does not produce a significant change in their relative comparison.
Source Right Ascension [rad] Declination [rad] $\ln\mathcal{B}^{\mathrm{GW+EM}}_{\mathrm{GW}}$
ZTF19abanrhr 3.35847 0.60781 N/A
BBHs $3.35^{+2.64}_{-0.14}$ $0.45^{+0.19}_{-1.34}$ 3.44
Echo $3.52^{+0.60}_{-0.28}$ $0.62^{+0.28}_{-1.38}$ 3.21